# Diagnosing Moral Reasoning Acquisition in Language Models: Pragmatics and Generalization

### v3

@@ Abstract
Guangliang Liu1      Lei Jiang2      Xitong Zhang1      Kristen Marie Johnson1 1Michigan State University     2University of Illinois Chicago {liuguan5,zhangxit,kristenj}@msu.edu    ljian43@uic.edu
1 Introduction and Related Works
@@
In cognitive science, Mahowald et al. (2024) suggest that while LLMs excel in formal language competence, they struggle with functional language competence—an essential requirement for acquiring moral reasoning capabilities. More fundamentally, Bender and Koller (2020) and other studies in BERTology Rogers et al. (2021), argue that Transformers cannot achieve true language acquisition, as it necessitates physical grounding and situated communicative intent Beuls and Van Eecke (2024), which extends beyond the distributional semantics captured by Transformers Harris (1954); Lenci et al. (2008); Boleda (2020). Previous studies Bonagiri et al. (2024); Zhang et al. (2023) demonstrate that LLMs do not have consistent moral or ethical orientations across various instances, which is contrary to the moral consistency principle Arvanitis and Kalliris (2020). Appendix A.1 contains additional related works and motivation pertaining to machine ethics.
@@
Specifically, we employ three fundamental tasks, Moral Foundations classification, rule of thumb generation, and ethical judgment prediction, as downstream evaluations of moral reasoning acquisition. We then compare their generalization characteristics with a representative semantics-driven task, sentiment analysis. Motivated by the distributional semantics hypothesis, we: (1) empirically show the generalization and convergence pitfalls of Moral Foundations classification; (2) propose the novel Representational Likelihood Algorithm (RLA) to statistically correlate representational similarity between seen and unseen moral pragmatics with the prediction likelihood of unseen morals; and (3) using RLA, perform mechanistic analysis of LLM performance gains for unseen situations.
@@ 2.2 Learning Paradigms
Notations. We denote the moral situation as $x_{s}$, the moral foundation as $y_{m}$, the RoT as $y_{r}$, and the judgment as $y_{j}$. Assuming an LLM $f$ is parameterized by $\theta$, RoT generation is formulated as $y_{r}=f_{\theta}(x_{s})$ and judgment prediction is represented as $y_{j}=f_{\theta}(x_{s})$.
Fine-tuning Strategies. Current learning paradigms of moral reasoning acquisition which aim to maximize conditional probabilities $\mathcal{P}_{\theta}(y_{r}|x_{s})$ and $\mathcal{P}_{\theta}(y_{j}|x_{s})$, typically apply a self-supervised fine-tuning or a reinforcement learning loss objective [^2] Please note the choice of objective loss function does not impact our conclusion.. Given the causal relationships among moral foundations, RoT, and judgment, previous studies often integrate them into a unified prediction task, such as $y_{r}=f_{\theta}(x_{s},y_{m})$ and $y_{j}=f_{\theta}(x_{s},y_{m},y_{r})$. During fine-tuning, the input for RoT generation can be $x_{s}$ with or without $y_{m}$, while the input for ethical judgment prediction can be $x_{s}$ with or without $y_{m}$ and/or $y_{r}$.
@@ 2.3 Pitfalls of Generalization
In this section, we use the Moral Foundations classification task as an example to illustrate its generalization pitfalls by comparing it to the semantics-level task of sentiment analysis. We argue that in moral classification tasks, there should be serious generalization issues since the classification model has to map semantically different situations into the same moral foundation label. A direct consequence is that an unseen situation is likely to be predicted correctly only if a semantically similar sample exists in the training set. This similarity requirement is much stricter than that for the sentiment analysis task.
@@
Observations for Classification Performance. Table 1 presents the classification performance on both the training and development set. Compared to the generalization behavior observed in SST (rightmost figure), the moral foundation classification tasks (first four figures) exhibit a significant performance gap between the training set and the development set. However, for MIC-RoT and SocialChem-RoT, because the training accuracy approaches 100% and converges after only several epochs, this suggests that task difficulty is not the primary cause of the observed generalization gap. The difference in classification performance between Situation and RoT stems from the fact that RoT is constructed based on typical moral foundations, inherently conveying information about the corresponding moral foundation. However, the generalization gap between the training set and development set for all moral foundation classification settings is apparent.
To further analyze the generalization pitfall in moral foundation classification, we examine the convergence behavior with respect to training dataset size. We use the curve of development accuracy in SST as a reference to highlight the convergence issue observed in moral foundation classification tasks.
@@ 3 Fine-tuning for Moral Reasoning Acquisition
moral-rot-judg .7628 .463 .345 .463 moral-rot-judg .761 .306 .000 .306
Table 3: Performance of Fine-tuned Mistral Model Across Various Fine-tuning Strategies for Each Benchmark, with the best strategy highlighted in bold. For both tasks, introducing more information, e.g., moral foundation, in the fine-tuning process would improve the performance. The moral-rot achieves the optimal performance for both SocialChem and MIC. The moral-judg and moral-judg are the best strategy for SocialChem and MIC respectively, in terms of the judgment prediction task. Additional results for Llama3 are availabe in Table 7.
Experimental Settings. We take Mistral-7B [^4] https://huggingface.co/mistralai/Mistral-7B-v0.1 and Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct as the backbone models and leverage the LoRA method to fine-tune them through a supervised fine-tuning loss. For each benchmark, we employ two fine-tuning strategies for RoT generation and four fine-tuning strategies for ethical judgment prediction. For RoT generation, we fine-tune LLMs: (1) to directly generate RoT according to the given situation (rot) and (2) first generate the moral foundation, then the RoT (moral-rot). For Judgment Prediction, we fine-tune LLMs to: (1) directly predict judgment (judg) , and (2) firstly generate the moral foundation and/or RoT then the judgment (moral-judgrot-judg and moral-rot-judg).
@@ 4.1 Representational Likelihood Algorithm
In this section, we present our novel method for identifying training samples that contribute to the prediction of a given test sample. We refer to these training samples as generalization-supportive samples [^7] In this paper, we use generalization-supportive and supportive interchangeably.. Our goal is to correlate representational similarity with LLM predictions, and then leverage this correlation to characterize the generalization mechanism of the considered morality acquisition tasks.
Assume that a fine-tuned LLM $f_{\theta}$ has been trained on the training set $\mathcal{D}_{\text{train}}$, where each sample is represented as $x=[x_{s},y_{m},y_{r},y_{j}]$, following the annotation introduced in Section 2.2. We denote training samples as $x\sim\mathcal{D}_{\text{train}}$ and test samples as $x^{\prime}\sim\mathcal{D}_{\text{test}}$. The hidden states of $f_{\theta}$ are denoted by $\mathcal{H}_{\theta}(\cdot)$, and the conditional likelihood of a given input and output is represented as $\mathcal{P}_{\theta}(\cdot|\cdot)$. Denote the cosine similarity function as $\text{cos}(\cdot)$.
@@
5:        $S^{t}$ = $\overbrace{\text{cos}(\mathcal{H}_{\theta}(x_{s}^{t}),\mathcal{H}_{\theta}(x_{s}^{{}^{\prime}}))}^{\text{{representational similarity}}}$cdot $\overbrace{\mathcal{P}_{\theta}(x^{t}_{j}|x^{t}_{s})}^{\textbf{likelihood}}$
6:        $\mathbf{d}[S^{t}]=\underbrace{\mathcal{P}_{\theta}(x^{t}_{j}|x^{{}^{\prime}}_{s})}_{\text{{prediction}}}$
@@
[…]
10:        $r$++
@@
Algorithm 1 presents our proposed Representational Likelihood Algorithm (RLA) by taking the judg fine-tuning strategy ($y_{j}=f_{\theta}(x_{s})$) as an instance. Specifically,
@@
    For each training sample $x^{t}$ in the sampled set $\mathcal{X}$, we calculate the similarity score $S^{t}$ which comprises the: (1) cosine similarity between two hidden states $\mathcal{H}_{\theta}(x_{s}^{t})$ and $\mathcal{H}_{\theta}(x_{s}^{{}^{\prime}})$ (line 5) measuring the representational similarity, and (2) likelihood, the conditional probability $\mathcal{P}_{\theta}(x^{t}_{j}|x^{t}_{s})$ measuring how good $f_{\theta}$ fits $x^{t}$ (line 5). With this design, only those training samples that have been fitted well by $f_{\theta}$ would be considered in the process of measuring representational similarity.
@@
    If $f_{\theta}$ becomes increasingly likely to assign $x_{t}$’s judgment $x^{t}_{j}$ to $x^{\prime}_{s}$ as their representational similarity increases, then we can correlate representational similarity and prediction (lines 8-10).
@@ 4.2 Interpretation of Generalization
Figure 4: Ratio of generalization-supportive training situations with the same underlying moral foudation as the test situation. Upper two subfigures are for SocialChem and the bottom sub-figures are for MIC.
@@
Figure 4 presents the results for this ratio. Interestingly, even for SST, which can be viewed as a binary classification task, only half of the supportive training samples share the same sentiment label as their corresponding test samples. For both RoT generation and ethical judgment prediction, the optimal fine-tuning strategies (moral-rot and moral-judg) align with the baseline fine-tuning strategies (rot and judg), except for moral-judg on the SocialChem benchmark. We believe this exception arises because the textual length of moral situations in SocialChem is relatively short, amplifying the influence of ethical judgment during fine-tuning. On the other hand, we calculate the average conditional likelihoods of the top-10 supportive training situations, and note the optimal fine-tuning strategy does help LLMs fit training samples. These observations suggest that LLMs consider moral situations and additional information together to generalize, but still operate primarily within the realm of semantics.
@@ A.5 Mechanistic Analysis to Fine-tuned Llama3
moral-rot-judg .752 .370 .248 .370 moral-rot-judg .762 .314 .000 .314
@@
Care Harm Demonstrates care, generosity, compassion, and empathy, while showing sensitivity to others’ suffering and upholding the principle of avoiding harm.
Fairness Cheating Encompasses fairness, justice, reciprocity, altruism, rights, autonomy, equality, proportionality, and the rejection of cheating.
Loyalty Betrayal Emphasizes group affiliation, solidarity, patriotism, and self-sacrifice, while prohibiting betrayal.
Authority Subversion Upholding social roles, respecting authority and traditions, valuing leadership, and prohibiting rebellion.
Purity (Sanctity) Degradation Reverence for the sacred, purity, religious principles guiding life, and prohibitions against violating the sacred.

### v5

@@ Abstract
Guangliang Liu1  Zimo Qi2  Xitong Zhang1  Lei Jiang3  Kristen Marie Johnson1 1Michigan State University  2Johns Hopkins University  3University of Illinois Chicago {liuguan5,zhangxit,kristenj}@msu.edu     zqi15@jh.edu    ljian43@uic.edu
1 Introduction
@@
In cognitive science, Mahowald et al. (2024) suggest that while LLMs excel in formal language competence, they struggle with functional language competence which is an essential requirement for acquiring moral reasoning capabilities. More fundamentally, Bender and Koller (2020) and other studies in BERTology Rogers et al. (2021), argue that Transformers cannot achieve true language acquisition, as it necessitates physical grounding and situated communicative intent Beuls and Van Eecke (2024), which extends beyond the distributional semantics captured by Transformers Harris (1954); Lenci et al. (2008); Boleda (2020). Previous studies Bonagiri et al. (2024); Zhang et al. (2023) demonstrate that LLMs do not have consistent moral or ethical orientations across various instances, which is contrary to the moral consistency principle Arvanitis and Kalliris (2020). Appendix A.1 contains additional related works and motivation pertaining to machine ethics.
@@
Specifically, we employ three fundamental tasks, Moral Foundations classification, rule of thumb generation, and ethical judgment prediction, as downstream evaluations of moral reasoning acquisition. We then compare their generalization characteristics with a representative semantics-driven task, sentiment analysis. Motivated by the distributional semantics theory, we: (1) empirically show the generalization and convergence pitfalls of Moral Foundations classification; (2) given the characteristic of autoregressive language models, propose a Representational Likelihood Algorithm (RLA) to statistically correlate representational similarity between seen and unseen situations with the prediction likelihood of unseen situations; and (3) using RLA, perform mechanistic analysis of LLM performance gains for unseen situations.
@@ 2.2 Learning Paradigms
Notations. We denote the moral situation as $m_{s}$, the moral foundation as $y_{m}$, the RoT as $y_{r}$, and the judgment as $y_{j}$. Assuming an LLM $f$ is parameterized by $\theta$, RoT generation is formulated as $y_{r}=f_{\theta}(m_{s})$ and judgment prediction is represented as $y_{j}=f_{\theta}(m_{s})$.
Fine-tuning Strategies. Current learning paradigms of moral reasoning acquisition which aim to maximize conditional probabilities $\mathcal{P}_{\theta}(y_{r}|m_{s})$ and $\mathcal{P}_{\theta}(y_{j}|m_{s})$, typically apply a self-supervised fine-tuning or a reinforcement learning loss objective [^2] Please note the choice of objective loss function does not impact our conclusion.. Given the causal relationships among moral foundations, RoT, and judgment, previous studies often integrate them into a unified prediction task, such as $y_{r}=f_{\theta}(m_{s},y_{m})$ and $y_{j}=f_{\theta}(m_{s},y_{m},y_{r})$. During fine-tuning, the input for RoT generation can be $m_{s}$ with or without $y_{m}$, while the input for ethical judgment prediction can be $m_{s}$ with or without $y_{m}$ and/or $y_{r}$.
@@ 2.3 Pitfalls of Generalization
In this section, we use the Moral Foundations classification task as an example to illustrate its generalization pitfalls by comparing it to the semantics-level task of sentiment analysis. We argue that in the moral foundations classification task, there should be serious generalization issues since the classification model has to map semantically different situations into the same moral foundation label. A direct consequence is that an unseen situation is likely to be predicted correctly only if a semantically similar sample exists in the training set. This similarity requirement is much stricter than that for the sentiment analysis task.
@@
Observations for Classification Performance. Figure 1 presents the classification performance on both the training and development set. Compared to the generalization behavior observed in SST (rightmost figure), the moral foundation classification tasks (first four figures) exhibit a significant performance gap between the training set and the development set. However, for MIC-RoT and SocialChem-RoT, because the training accuracy approaches 100% and converges after only several epochs, this suggests that task difficulty is not the primary cause of the observed generalization gap. The difference in classification performance between Situation and RoT stems from the fact that RoT is constructed based on typical moral foundations, inherently conveying information about the corresponding moral foundation. However, the generalization gap between the training set and development set for all moral foundation classification settings is apparent. To further analyze the generalization pitfall in moral foundation classification, we examine the convergence behavior with respect to training dataset size. We use the curve of development accuracy in SST as a reference to highlight the convergence issue observed in moral foundation classification tasks.
@@ 3 Fine-tuning for Moral Reasoning Acquisition
moral-rot
-judg
.7628 .463 .345 .463
moral-rot
-judg
.761 .306 .000 .306
Table 3: Performance of Fine-tuned Mistral Model Across Various Fine-tuning Strategies for Each Benchmark, with the best strategy highlighted in bold. For both tasks, introducing more information, e.g., moral foundation, in the fine-tuning process would improve the performance. The moral-rot achieves the optimal performance for both SocialChem and MIC. The moral-judg is the best strategy for both SocialChem and MIC, in terms of the judgment prediction task. Additional results for Llama3 are availabe in Table 7.
Experimental Settings. We take Mistral-7B [^4] https://huggingface.co/mistralai/Mistral-7B-v0.1 and Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct as the backbone models and leverage the LoRA method to fine-tune them through a supervised fine-tuning loss. For each benchmark, we employ two fine-tuning strategies for RoT generation and four fine-tuning strategies for ethical judgment prediction. For RoT generation, we fine-tune LLMs: (1) to directly generate RoT according to the given situation (rot) and (2) first generate the moral foundation, then the RoT (moral-rot). For Judgment Prediction, we fine-tune LLMs to: (1) directly predict judgment (judg) , and (2) firstly generate the moral foundation and/or RoT then the judgment (moral-judg, rot-judg and moral-rot-judg).
@@ 4.1 Representational Likelihood Algorithm
Motivated by the representation similarity hypothesis in domain generalization Ben-David et al. (2006), we present our method for identifying training samples that contribute to the prediction of a given test sample. We refer to these training samples as generalization-supportive samples [^7] In this paper, we use generalization-supportive and supportive interchangeably.. Our goal is to correlate representational similarity with LLM predictions, and then leverage this correlation to characterize the generalization mechanism of the considered morality acquisition tasks.
Assume that a fine-tuned LLM $f_{\theta}$ has been trained on the training set $\mathcal{D}_{\text{train}}$, where each sample is represented as $x=[m_{s},y_{m},y_{r},y_{j}]$, following the annotation introduced in Section 2.2. We denote training samples as $x\sim\mathcal{D}_{\text{train}}$ and test samples as $x^{\prime}\sim\mathcal{D}_{\text{test}}$. The hidden states of $f_{\theta}$ are denoted by $\mathcal{H}_{\theta}(\cdot)$, and the conditional likelihood of a given input and output is represented as $\mathcal{P}_{\theta}(\cdot|\cdot)$. Denote the cosine similarity function as $\text{cos}(\cdot)$.
@@
5:   $S^{t}$ = $\overbrace{\text{cos}(\mathcal{H}_{\theta}(m_{s}^{t}),\mathcal{H}_{\theta}(m_{s}^{{}^{\prime}}))}^{\text{{representational similarity}}}$ $\cdot$ $\overbrace{\mathcal{P}_{\theta}(y^{t}_{j}|m^{t}_{s})}^{\textbf{likelihood}}$
6:   $\mathbf{d}[S^{t}]=\underbrace{\mathcal{P}_{\theta}(y^{t}_{j}|m^{{}^{\prime}}_{s})}_{\text{{prediction}}}$
@@
9:  if MEAN($\mathcal{V}$$\frac{\mathcal{N}}{2}]$) << MEAN($\mathcal{V}[\frac{\mathcal{N}}{2}:]$) then
10:   rr++
@@
Algorithm 1 presents our proposed Representational Likelihood Algorithm (RLA) by taking the judg fine-tuning strategy ($y_{j}=f_{\theta}(m_{s})$) as an instance. Specifically,
@@
    For each training sample $x^{t}$ in the sampled set $\mathcal{X}$, we calculate the similarity score $S^{t}$ which comprises the: (1) cosine similarity between two hidden states $\mathcal{H}_{\theta}(m_{s}^{t})$ and $\mathcal{H}_{\theta}(m_{s}^{{}^{\prime}})$ (line 5) measuring the representational similarity, and (2) likelihood, the conditional probability $\mathcal{P}_{\theta}(y^{t}_{j}|m^{t}_{s})$ measuring how good $f_{\theta}$ fits $x^{t}$ (line 5). With this design, only those training samples that have been fitted well by $f_{\theta}$ would be considered in the process of measuring representational similarity.
@@
    If $f_{\theta}$ becomes increasingly likely to assign $m_{s}^{t}$’s judgment $y^{t}_{j}$ to $m^{\prime}_{s}$ as their representational similarity increases, then we can correlate representational similarity and prediction (lines 8-10).
@@ 4.2 Interpretation of Generalization
Figure 4: Ratio of generalization-supportive training situations with the same underlying moral foudation as the test situation. Upper two subfigures are for SocialChem and the bottom sub-figures are for MIC. Top-50 situations are available in Appendix A.6.
@@
Figure 4 presents the results for this ratio. Interestingly, even for SST, which can be viewed as a binary classification task, only half of the supportive training samples share the same sentiment label as their corresponding test samples. For both RoT generation and ethical judgment prediction, the optimal fine-tuning strategies (moral-rot and moral-judg) align with the baseline fine-tuning strategies (rot and judg), except for moral-judg on the SocialChem benchmark. We believe this exception arises because the textual length of moral situations in SocialChem is relatively short, amplifying the influence of ethical judgment during fine-tuning.
On the other hand, Table 5 reports the average conditional likelihoods of the top-10 supportive training situations, and note the optimal fine-tuning strategy does help LLMs fit training samples. These observations suggest that LLMs consider moral situations and additional information together to generalize, but still operate primarily within the realm of semantics.
@@
5 Discussion
Generalization remains a significant challenge in the acquisition of moral reasoning, and no optimal solution has yet been identified. Recently, Jiang et al. (2025) proposed a hybrid approach that combines bottom-up and top-down methods. However, their method still relies on a substantial number of training samples. Bergen et al. (2016) demonstrated that pragmatic reasoning can be approximated through semantic inferences, highlighting a linguistic foundation for this connection. Nevertheless, how to formally structure a semantic inference framework for moral reasoning remains an open question. One promising direction is to ground such a framework in the human moral decision-making process. Kumar and Jurgens (2025) introduced the first benchmark in the NLP community focused on how humans make moral decisions. Their benchmark is based on an intuitionist model: participants are first asked to make a moral judgment and then provide an explanation for their decision. This type of annotation presents challenges for LLMs, as human explanations are expressed in free-text form and often lack enough situated semantic information Sap et al. (2022). Despite these difficulties, the benchmark offers a valuable opportunity for exploring methods that aim to derive semantic inferences from human rationales, potentially bridging the gap in pragmatic reasoning for morality.
Sap et al. (2019) propose a social bias inference framework to identify social implications, but their approach still relies heavily on semantics-level signals, as the social implications in their work are primarily associated with explicit toxicity. Chen and Wang (2025) introduce a pragmatic inference framework targeting implicit toxicity, where metaphorical expressions of toxicity pose particular challenges for LLMs. They provide empirical evidence that their framework enhances the detection of implicit toxicity in off-the-shelf LLMs, though it still depends on the powerful distributional semantics captured by these models. We argue that the study of moral reasoning acquisition should focus on smaller LLMs with 3B–8B parameters, as these models are less powerful than very large ones yet still exhibit satisfactory instruction-following capabilities. The powerful distributional semantics of large models make it difficult to disentangle whether performance gains stem from the proposed pragmatic inference framework or from learned statistical correlations between situations and moral reasoning objectives.
@@ A.5 Mechanistic Analysis to Fine-tuned Llama3
moral-rot
-judg
.752 .370 .248 .370 moral-rot -judg .762 .314 .000 .314
@@
Care
Harm
Demonstrates care, generosity, compassion, and empathy,
while showing sensitivity to others’ suffering and upholding the principle of avoiding harm.
Fairness
Cheating
Encompasses fairness, justice, reciprocity, altruism, rights,
autonomy, equality, proportionality, and the rejection of cheating.
Loyalty
Betrayal
Emphasizes group affiliation, solidarity, patriotism,
and self-sacrifice, while prohibiting betrayal.
Authority
Subversion
Upholding social roles, respecting authority and
traditions, valuing leadership, and prohibiting rebellion.
Purity (Sanctity)
Degradation
Reverence for the sacred, purity, religious principles guiding life,
and prohibitions against violating the sacred.
@@
A.6 Top-50
In Figure 4, we show only 10 generalization-supportive samples. Here, we demonstrate that the characteristics of all top-50 generalization-supportive training samples are closely aligned with those of the top 10 reported in that figure.
Mistral-SocialChem-RoT:[0.138, 0.16, 0.221, 0.193, 0.189, 0.185, 0.21, 0.193, 0.17, 0.178, 0.18, 0.176, 0.181, 0.191, 0.187, 0.157, 0.161, 0.145, 0.163, 0.133, 0.15, 0.181, 0.152, 0.162, 0.18, 0.163, 0.173, 0.16, 0.158, 0.186, 0.176, 0.178, 0.17, 0.185, 0.171, 0.169, 0.165, 0.194, 0.191, 0.173, 0.19, 0.173, 0.188, 0.192, 0.188, 0.195, 0.189, 0.19, 0.195, 0.17] with mean value of 0.17636
Mistral-Socialchem-MoralRoT: [0.051, 0.243, 0.238, 0.214, 0.079, 0.245, 0.244, 0.244, 0.241, 0.216, 0.072, 0.133, 0.204, 0.276, 0.137, 0.179, 0.178, 0.115, 0.049, 0.151, 0.152, 0.16, 0.089, 0.048, 0.186, 0.141, 0.126, 0.137, 0.146, 0.047, 0.045, 0.041, 0.122, 0.156, 0.143, 0.084, 0.237, 0.232, 0.135, 0.099, 0.09, 0.207, 0.371, 0.169, 0.23, 0.127, 0.093, 0.199, 0.164, 0.163] with mean of 0.15696